A store sells cashews for $4.40/lb and peanuts for $1.20/lb. How many lb of each can be bought to get exactly 3lb for $6. If possible please solve using system of equations.
Answer by subudear(62)

This can be solved using system of equations.
Let buy x lb of cashew and y lb of peanuts.
x + y = 31 .....(1)
Cost of cashew and peanuts will be
4.4x + 1.2y = 6 ......(1)
Solving these equation will provide the solution.
lets replace y = 31 - x from equation (1) into equation (2)
4.4x + 1.2 (31 - x) = 6
4.4x + 37.2 - 1.2x = 6
3.2x = 31.2
x = 9.75
substituting the value sin equation (1) will give
y = 31 - 9.75 = 21.25
Cashews weight will be 9.75 lb
peanuts weight will be 21.25 lb